The Ohio Supercomputer Center (OSC) is experiencing an email delivery problem with several types of messages from MyOSC. 

Known issues

Unresolved known issues

Known issue with an Unresolved Resolution state is an active problem under investigation; a temporary workaround may be available.

Resolved known issues

A known issue with a Resolved (workaround) Resolution state is an ongoing problem; a permanent workaround is available which may include using different software or hardware.

A known issue with Resolved Resolution state has been corrected.

Known Issues

Title Category Resolution Description Posted Updated
Poor network performance on some filesystems filesystem Resolved

We are experiencing some network performance issues on a cluster of servers involved with providing GPFS and some project filesystems. GPFS appears to be functioning acceptably, but proj01, proj02... Read more

10 years 9 months ago 10 years 9 months ago
Network card re-seat Network Resolved

At 8AM on Tuesday, July 9th 2013, we will be re-seating a network card in a switch at our operations center. It is possible that a brief (~10 minute) outage may occur. Jobs will pause for the... Read more

10 years 9 months ago 10 years 9 months ago
6/4/13 Scheduled Downtime Outage Resolved

HPC systems are currently offline for scheduled maintenance. See osc.edu/n for more information.

10 years 10 months ago 10 years 10 months ago
Brief interruption of services for some users filesystem Resolved

Today, May 14 2013, at 12:45PM we will be temporarily removing one of the home directory servers from service to address some reliability issues. Users with home... Read more

10 years 11 months ago 10 years 11 months ago
Backups of /nfs/gpfs Backups Resolved

Changes to files on /nfs/gpfs may not be backed up during the following evening's backup, as would normally be expected. The backup software is attempting to recreate a full backup of the... Read more

11 years 2 days ago 10 years 11 months ago
Brief interruption of batch services on 4/17 Batch Resolved

On April 17th 2013, at roughly 2PM, we will be rebooting the batch server on the Oakley cluster. Running jobs will not be affected, but there will be a brief disruption in scheduling, as well as... Read more

11 years 1 week ago 11 years 1 week ago
16 core nodes on Glenn temporarily unavailable Operations Resolved

This issue has been resolved. The 16-core nodes are online.

---------------------------------------------------------------------------------------

16 core nodes on Glenn are currently... Read more

11 years 1 month ago 11 years 3 weeks ago
2/26 Downtime Difficulties Operations Resolved

All systems should be functioning normally. Please report any remaining issues to OSC Help.

----
A number of systems are still experiencing problems after yesterday's... Read more

11 years 1 month ago 11 years 1 month ago

Pages